A client of the ABC investment company wanted to invest $20,000. She was advised to invest in a mutual bond that earns 3.5% annual simple interest and a stock fund that earns 9.3% annual simple interest. How much should be invested in each account so that both accounts earn the same annual interest?
Answer by Nate(3500)

I = PRT
Interest Yields Principle * Rate * Time
mutual bond: I = 0.035*P*T
stock fund: I = 0.093*P*T
The interests are the same:
0.035*P*T = 0.093*P*T
0.035*P = 0.093*P
~> set m = amount in mutual bonds ~> set 20,000 - m = amount in stock fund
0.035(m) = 0.093(20,000 - m)
0.035m = 1,860 - 0.093m
0.128m = 1,860
m = 14,531.25
20,000 = 5,468.75
Mutual Bonds: $14,531.25
Stock Funds: $5,468.75